OpenStudy (anonymous):

Find a & b if squareroot of 13 - squareroot of 11 are divided by squareroot of 13 + squareroot of 11 and is equal to a - b times squareroot of 143

OpenStudy (dumbcow):

\[(a-b)\sqrt{143} = \frac{\sqrt{13}-\sqrt{11}}{\sqrt{13}+\sqrt{11}}\]

OpenStudy (angela210793):

yea

OpenStudy (angela210793):

then (a-b)sqr143=24-2sqr143

OpenStudy (anonymous):

two unknows, u must have two eqs

OpenStudy (dumbcow):

thats what im thinking

OpenStudy (angela210793):

(a-b)=(24-2sqr143)/sqr143 can't go any further

OpenStudy (anonymous):

\[(a-b)\sqrt{143}={{2} \over{(\sqrt{13}-\sqrt{11})^2}}\]

OpenStudy (dumbcow):

\[a-b = \frac{12-\sqrt{143}}{\sqrt{143}}\]

OpenStudy (anonymous):

\[2=(a-b)\sqrt{143}(24-\sqrt{143})\]

OpenStudy (angela210793):

yea but there must b smth else to find a and b

OpenStudy (anonymous):

yea another relation in a n b

OpenStudy (dumbcow):

technically there are an infinite number of solutions pick a value for b then assign appropriate value for a

OpenStudy (anonymous):

(a-b) = something so a = something + b and b = a - somthing